Find all values of x where the function f(x)=-4x^2/e^-4x has a relative maximum. If there are no such values of x, submit an empty answer.


1
Expert's answer
2021-01-12T14:51:09-0500

To do this we will first calculate the derivative :

"\frac{d}{dx}(\frac{-4x^2}{e^{-4x}})=\frac{d}{dx}(-4x^2e^{4x}) = -8xe^{4x}-16x^2e^{4x}=-8x(1+2x)e^{4x}"

The roots are "x_1 = 0, x_2 = -\frac{1}{2}" .

Now we will analize the sign of "f'(x)" near these points :

At "x_1=0, f'" changes it's sign from "+" to "-" (exponent is always positive, "(1+2x)" remains positive around "x_1"), thus "x_1" is a local maximum (and we can even see that it is a global maximum).

At "x_2=-\frac{1}{2}, f'" changes it's sign from "-" to "+" ("x" remains negative around "x_2"), so "x_2" is a local minimum and therefore is not a local maximum.